Session: Mar 2008 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Sat Mar 1 11:06:15 2008) Success :
Windy at the end (and rain). Otherwise no known problems.

Urumqi (Hua Zhang at Sat Mar 1 12:33:22 2008) Success :

Torun (Kaz Borkowski at Mon Mar 3 07:10:01 2008) Success :
No known problems.

Medicina (Andrea Orlati at Mon Mar 3 07:39:15 2008) Minor failures : RFI

Westerbork (Tony Foley at Mon Mar 3 09:47:26 2008) Minor failures :
Wb used in phased array mode. One of the disk packs has the wrong VSN; in the log file it says OSOD-057 but the external label is JIVE-020

Effelsberg (Uwe Bach at Mon Mar 3 12:18:04 2008) Minor failures : WIND
Strong wind! Observations were stopped at about 22:50 UT

Jodrell_Bank (Alastair Gunn at Mon Mar 3 15:49:14 2008) Failure : WIND
The Lovell Telescope was used (Jb1). Data recorded but telescope was offline for entire experiment due to high winds.

Cambridge (Alastair Gunn at Mon Mar 3 15:49:49 2008) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 4 17:37:30 2008) Failure : RECO
Error in pack change, not noticed by the operator.

Noto (P. Cassaro at Tue Mar 4 17:41:03 2008) Failure : RECO
The experiment was recorded until 18:00 UT.

Hartebeesthoek (Jonathan Quick at Sat Mar 8 09:09:39 2008) Success : RECE
Receiver still cooling down until 17:45 UT but seems to have worked OK